HIGH RESOLUTION INFRARED SPECTROSCOPY ON LINEAR CLUSTER IONS

\underline{H. VERBRAAK}, H. LINNARTZ, Sackler Laboratory for Astrophysics, Leiden Observatory, Leiden University, P.O. Box 9513, NL-2300 RA Leiden, The Netherlands; M. SNELS, Istituto Scienze dell’ Atmosfera e del Clima, Via Fosso del Cavaliere 100, Italy; D. VERDES, Institute for Physical Chemistry, University of Zürich, Winterhurerstrasse 190, CH 8057 Zürich, Switzerland; and P. BOTSCHWINA, Institut für Physikalische Chemie, Tammannstrasse 6, D-37077 Göttingen, Germany.

A general applicable technique for high resolution infrared direct absorption spectroscopy of ionic complexes is presented. The setup uses the output of tunable diode lasers to sample a planar plasma based upon electron impact ionization of a gas mixture that is expanded supersonically through a long and narrow slit. A rich variety of ionic complexes and cluster ions is formed in the expansion^{a} with abundances high enough to be detectable in a production modulation scheme. The performance of the setup is demonstrated on recently obtained spectra of the ν_{1} fundamental of the Ar···DN_{2}^{+} complex that coincides with the ν_{2}+4ν_{s} band resulting in a beautiful example of a strong Fermi interaction^{b}.

References a H. Verbraak, D. Verdes, H. Linnartz, Int. J. Mass Spectrom. to be published (doi:10.1016/j.ijms.2007.02.042) b H. Verbraak, M. Snels, P. Botschwina, H. Linnartz, J. Chem. Phys. \underline{124}, 224315 (2006) c H. Verbraak, J. N. P. van Stralen, J. Bouwman, J. S. de Klerk, D. Verdes, H. Linnartz, F. M. Bickelhaupt, J. Chem. Phys. \underline{123}, 144350 (2005)